OpenStudy (anonymous):

Derive y = lnx / (x^1/6)

OpenStudy (anonymous):

find Derivative ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u=lnx =>du=1/x dv=x^(1/6) =>v=6/5x^(5/6) \[\int\limits_{}^{}=u*v-\int\limits_{}^{}vdu\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

you can write it as \(\huge y=(x^{-\frac{1}{6}})\times lnx \) and use the product rule.... much simpler...

OpenStudy (anonymous):

I confusion antiderivative

OpenStudy (anonymous):

(uv)'=u'v+uv'

OpenStudy (anonymous):

\(\large y'=[x^{-\frac{1}{6}}]'lnx+[lnx]' x^{-\frac{1}{6}} \)

OpenStudy (anonymous):

can that be simplified any further?